write y in terms of x?

(-x/3)+(y/2)=(5/8)
I dont understand this? Are we solving for y or x?

Answer by alicealc(281)

it means that we have to make it y = .... x

(-x/3) + y/2 = 5/8
y/2 = 5/8 + x/3
multiply each side by 2:
y = 10/8 + 2x/3
y = 5/4 + 2x/3